# Cutting Through Privacy: A Hyperplane-Based Data Reconstruction Attack in Federated Learning

### v1

@@ 4.3 Controlling the Gradients
In this way, the partial derivatives become independent of $b^{(1)}_{i}$ and $\alpha_{j}$ can be artificially set to a non-negligible value without knowledge of the $y_{j}$, as required by Assumption 4.1. While this reasoning holds for any choice of the matrix $\mathbf{W}^{(2)}$, we require ${\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i}}\neq 0$ for the input ${\mathbf{x}}_{j}$ to have a non-zero coefficient in the aggregate gradient. Furthermore, for the efficient parallel search described in the next section, we want ${\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i_{1}}}={\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i_{2}}}$ across all neurons, which requires the rows of $\mathbf{W}^{(2)}$ to be identical. These constraints are easily satisfied since the attacker has control over all model parameters.
@@ 4.4 Parallel Search
The initial search space for the bias values, $[l_{0},u_{0}]$, is determined based on the range of possible input feature values. In the first communication round, this interval is divided into $N$ equal-length subintervals. The procedure UpdateSearchState in Algorithm 2 identifies which subintervals require further exploration. Subintervals containing no inputs, as well as those whose size becomes smaller than the threshold $\varepsilon$, are discarded. At any round, $\mathcal{I}$ represents the set of remaining subintervals. The attacker tests $N$ new bias values within these subintervals, distributing the values roughly evenly among them. This procedure, called UpdateHyperplanes, is detailed in Algorithm 3 in App. E.
The number of rounds required for full recovery of the batch grows logarithmically with the batch size $n$ (App. D).
@@
Input: the set of search intervals $\mathcal{I}$, the malicious model parameters $\theta=\{\mathbf{W}^{(1)},\mathbf{W}^{(2)},{\mathbf{b}}^{(1)},{\mathbf{b}}^{(2)}\}$, the set of attack rounds $\mathcal{T}=\{1,...,T\}$
1:  $\mathcal{G}\leftarrow\{\}$
2:  $\mathcal{I}\leftarrow[l_{0},u_{0}]$
3:  for $t\in\{1,...,T\}$  do
4:     ${\mathbf{b}}^{(1)}$ $\leftarrow\mathrm{UpdateHyperplanes}(\mathcal{I},{\mathbf{b}}^{(1)})$
5:     server sends malicious parameters $\theta$ to the client
6:     client sends gradient updates $\frac{\partial{\mathcal{L}_{j}}}{\partial{\mathbf{W}^{(1)}_{i}}}$ and $\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}}$, for $i=\{1,...,N\}$
7:     server computes $\mathbf{g}_{i}$ using (3)
8:     $\mathcal{G}\leftarrow\mathcal{G}\cup(\mathbf{g}_{i},\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}},b_{i}^{(1)})$, for $i=\{1,...,N\}$
9:     $\mathcal{G},\mathcal{I}\leftarrow\mathrm{UpdateSearchState}(\mathcal{G})$
10:  end for
11:  Reconstruct the input from $\mathcal{G}$ using (8)
@@
  {}$\mathcal{I}\leftarrow\{\},\;\mathcal{G}_{\mathrm{new}}\leftarrow\{\}$
  sort $\mathcal{G}$ by ascending value of the bias values
  for $i=2,\dots,|\mathcal{G}|$ do
     if $\mathbf{g}_{i}\neq\mathbf{g}_{i-1}$ and $b_{i}-b_{i-1}\geq\varepsilon$  then
        $\mathcal{G}_{\mathrm{new}}\leftarrow\mathcal{G}_{\mathrm{new}}\cup\left(\mathbf{g}_{i},\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}},b_{i}^{(1)}\right)$
        $\mathcal{I}\leftarrow\mathcal{I}\cup[b_{i-1}^{(1)},b_{i}^{(1)}]$
     end if
  end for
  Return $\mathcal{G}_{\mathrm{new}},\mathcal{I}$
@@ 5 Experiments
We conduct experiments to demonstrate the efficacy of our proposed attack on both image and tabular data. Our evaluation encompasses two distinct datasets, each representing a different data modality. For image classification tasks, we use the ImageNet ILSVRC 2012 dataset [Russakovsky et al., 2015], that comprises a total of 1,000 classes. As preprocessing steps, all images are rescaled to the $[0,1]$ range and resized to $224\times 224$ pixels. To evaluate performance on tabular data, we test the Human Activity Recognition Using Smartphones dataset [Anguita et al., 2013] (HARUS), which includes recordings of 30 subjects performing various activities while carrying a waist-mounted smartphone equipped with embedded sensors. The dataset comprises 561 features and 6 classification labels, representing the activity performed by each user. All features are scaled between $-1$ and $1$. Each client trains a FC-NN with two layers. Unless stated otherwise, the first layer consists of $N=1,000$ neurons. For all the experiments, we assume that training is performed using FedSGD with full-batch updates. An image is considered reconstructed if it has SSIM $\geq 0.99$ relative to the true image (Figures 7 and 8 provide examples of reconstructed images). For tabular data, we consider an input to be perfectly reconstructed if the L2-norm difference between the true and recovered data point is $<0.1$.
We compare our method to the Curious Abandon Honesty (CAH) attack [Boenisch et al., 2023], allowing, for a fair evaluation, their attack to be performed across multiple communication rounds. Each experiment is repeated three times with different random seeds to ensure consistency and reliability of the results. Details on the configuration of both attacks can be found in App. F.
@@ 5.1 Experimental Results
Figure 3 illustrates the performance difference between our attack and CAH on the HARUS dataset. Two key trends emerge from our results. First, the performance gap increases with both the number of communication rounds and the batch size. This is clearly visible in the figure, with the largest performance difference ($97.75$ percentage points) observed at 50 rounds and a batch size of 4,096. Second, different regions of the heatmap reveal the relative strengths of each approach. The top-left region (small batches, many rounds) suggests CAH’s effectiveness on smaller batches; the bottom row highlights its advantage in one-shot attacks. However, this advantage relies on their trap weights initialization, which requires the attacker to have access to some representative data to tune some hyperparameters (see App. F). This relative advantage disappears as the batch size increases. The top-right region (large batches, many rounds) highlights our attack’s ability to scale, demonstrating its potential for larger batches. This supports our claim that a sufficient number of rounds enables recovery for arbitrary large batches.
We also evaluate the effect of the number of neurons $N$ in the fully connected attack layer and its relationship with attack accuracy, see Table 1. While we observe that the percentage of reconstructed data increases with a larger number of neurons for both attacks, our attack outperforms the baseline in all but one combination of rounds and neuron counts. In many cases, our attack reconstructs nearly 100 times more inputs than CAH. We highlight a key distinction: the baseline reconstructs around 3.6% of the input after 50 rounds with 2000 neurons, i.e., using 100,000 different random hyperplanes. On the other hand, our attack manages to reconstruct 38% of the batch after only 10 rounds with only 500 neurons, i.e., using only 5,000 different hyperplanes. In comparison, our attack reconstructs more than 10 times as many input points using 20 times fewer hyperplanes than the baseline, thanks to its efficient search procedure for correctly positioning the hyperplanes that isolate the inputs.
We provide additional experimental results on HARUS, ImageNet, and CIFAR-10 [Krizhevsky, 2009] in App. G, including tests on a CNN architecture. These results confirm the same qualitative trends observed in our main results, further validating our conclusions.
@@ 6 Discussion
Secure aggregation [Bonawitz et al., 2017] is often proposed as a defense mechanism against privacy attacks, as it aggregates updates from multiple clients without revealing individual client updates to the server. However, the server still obtains the average of the gradients computed over the entire dataset. With a sufficient number of communication rounds, our attack could still recover all underlying data, though not necessarily associate it with a specific client.
Therefore, the only robust defense against our attack that we are aware of would involve implementing local differential privacy (LDP) [Dwork, 2006, Bhowmick et al., 2019] on the client side. By adding noise to updates, LDP alters the coefficients associated to each input in (3), thus hindering precise localization and subsequent reconstruction of the data points, at the cost of decreased model utility.
@@ 7 Future Works and Conclusion
Extending this attack to settings where clients perform multiple local steps presents an interesting challenge. One potential approach could involve leveraging the server’s control over client training hyperparameters by setting an exceptionally low learning rate. This would ensure that local updates to the input layer would not significantly alter the orientation of the hyperplanes, thereby preserving the feasibility of the reconstruction. A similar result can be obtained by modifying the activation function so that updates are clipped. However, handling potential updates across multiple mini-batches remains an open question.
Exploring the applicability of the attack to other machine learning tasks and architectures (beyond those discussed in Sec. 2.2.3) represents another promising direction for future work.
@@ A Hyperplane-Based Data Reconstruction Attack in Federated Learning
By definition of the method of attack, a point can only be reconstructed using (2) if it is linearly separated by injecting malicious weights. Thus, the number of recovered points is upper-bounded by the number of linearly separable points in the batch $\mathcal{B}_{{\mathbf{x}}}$, which corresponds to the number of vertices on its convex hull. This value is bounded by
@@
    $O(n^{(d-1)/(d+1)})$ if $\mathcal{B}_{{\mathbf{x}}}$ is sampled uniformly at random from a unit ball Raynaud [1970], or
@@
    $O(\log^{d-1}n)$ if $\mathcal{B}_{{\mathbf{x}}}$ is sampled uniformly at random from the unit hypercube Bentley et al. [1978], or
@@
    $O(\log^{(d-1)/2}n)$ if $\mathcal{B}_{{\mathbf{x}}}$ is sampled from a normal distribution with covariance matrix $\mathbf{I}_{d}$ Raynaud [1970].∎
@@ Appendix B Partial derivative with respect to the bias
$\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}}=\frac{\partial{\mathcal{L}_{j}}}{\partial{z^{(1)}_{i}}}=\frac{\partial{\mathcal{L}_{j}}}{\partial{{\mathbf{z}}^{(2)}}}\frac{\partial{{\mathbf{z}}^{(2)}}}{\partial{z_{i}^{(1)}}},$ (11)

### v2

@@ 4.3 Controlling the Gradients
In this way, the partial derivatives become independent of $b^{(1)}_{i}$ and $\alpha_{j}$ can be artificially set to a non-negligible value without knowledge of the $y_{j}$.While this reasoning holds for any choice of the matrix $\mathbf{W}^{(2)}$, we require ${\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i}}\neq 0$ for the input ${\mathbf{x}}_{j}$ to have a non-zero coefficient in the aggregate gradient. Furthermore, for the efficient parallel search described in the next section, we want ${\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i_{1}}}={\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i_{2}}}$ across all neurons, which requires the columns of $\mathbf{W}^{(2)}$ to be identical. These constraints are easily satisfied since the attacker has control over all model parameters.
@@ 4.4 Parallel Search
The initial search space for the bias values, $l_{1},u_{1}]$, is determined based on the range of possible input feature values. In the first communication round, this interval is divided into $N$ equal-length subintervals. The procedure UpdateSearchState in Algorithm 2 identifies which subintervals require further exploration. Subintervals containing no inputs, as well as those whose size becomes smaller than the threshold $\varepsilon$, are discarded.
At each round, $\mathcal{I}$ represents the set of remaining subintervals. The attacker tests $N$ new bias values within these subintervals, distributing the values roughly evenly among them. More precisely, if $N$ cannot be evenly divided among the subintervals, the extra bias values are assigned to the longer intervals. This procedure, called UpdateHyperplanes, is detailed in Algorithm. 3.
@@
Input: $l_{1},u_{1}]$ initial search space of bias values, the set of attack rounds $\mathcal{T}=\{1,...,T\}$
[…]
2:{[$\mathcal{I}\leftarrow\{[l_{1},u_{1}]\}$
3: Draw a vector $\mathbf{w}\in\mathbb{R}^{d}$ and a vector $\mathbf{v}\in\mathbb{R}^{C}$, set each row of $\mathbf{W}^{(1)}$ to $\mathbf{w}$ and each column of $\mathbf{W}^{(2)}$ to $\mathbf{v}$
4: Set all entries of ${\mathbf{b}}^{(2)}$ to the same large value
5:for $t\in\{1,...,T\}$  do
6:  ${\mathbf{b}}^{(1)}$ $\leftarrow\mathrm{UpdateHyperplanes}(\mathcal{I})$
7:  server sends malicious parameters $\theta$ to the client
8:  server receives gradient updates $\frac{\partial{\mathcal{L}}}{\partial{\mathbf{W}^{(1)}}}$ and $\frac{\partial{\mathcal{L}}}{\partial{{\mathbf{b}}^{(1)}}}$
9:  server computes $\mathbf{g}_{i}$ using (5), $\forall i\in\{1,...,N\}$
10:  $\mathcal{G}\leftarrow\mathcal{G}\cup\{(\mathbf{g}_{i},\frac{\partial{\mathcal{L}}}{\partial{b^{(1)}_{i}}},b_{i}^{(1)}),\forall i=\{1,...,N\}\}$
11:  $\mathcal{G},\mathcal{I}\leftarrow\mathrm{UpdateSearchState}(\mathcal{G})$
12: Reconstruct the inputs from $\mathcal{G}$ using (8)
@@
[…]
2: sort $\mathcal{G}$ by ascending value of the bias values
3:for $i=2,\dots,|\mathcal{G}|$ do
4:  if $\mathbf{g}_{i}\neq\mathbf{g}_{i-1}$ and $b_{i}-b_{i-1}\geq\varepsilon$  then
5:   $\mathcal{G}_{\mathrm{new}}\leftarrow\mathcal{G}_{\mathrm{new}}\cup\{\left(\mathbf{g}_{i},\frac{\partial{\mathcal{L}}}{\partial{b^{(1)}_{i}}},b_{i}^{(1)}\right)\}$
6:   $\mathcal{I}\leftarrow\mathcal{I}\cup\{[b_{i-1}^{(1)},b_{i}^{(1)}]\}$
7: Return $\mathcal{G}_{\mathrm{new}},\mathcal{I}$
Algorithm 3 UpdateHyperplanes
Input: set of search intervals $\mathcal{I}=\{[l_{1},u_{1}],[l_{2},u_{2}],\ldots,[l_{k},u_{k}],\ldots\}$
[…]
2:if $M=1$ then
3:  ${\mathbf{b}}^{(1)}\leftarrow\big{\{}l_{1}+i\frac{u_{1}-l_{1}}{N+1}|i=1,\dots,N\big{\}}$
4:else
5:  sort $\mathcal{I}$ by descending length $(u_{k}-l_{k})$
7:  for k=1,,Mk=1,...,M do
8:   if $k<r$ then
10:    $j\leftarrow j+q+1$
11:   else
13:    $j\leftarrow j+q$
14: Return ${\mathbf{b}}^{(1)}$
4.5 Complexity Analysis
To evaluate the time complexity of the attack (Alg. 1), we decompose the operations performed by the server during each attack round, and subsequently analyze the total computational cost.
At each round, the server first prepares the malicious hyperplanes for the targeted client (line 6). This procedure requires $\mathcal{O}(n\log n)$ operations, primarily due to the sorting operation on the set of subintervals $\mathcal{I}$ of size at most $n$ (line 5 in Alg. 3). [^2] As we can see from Alg. 2, $\mathcal{I}$ and $\mathcal{G}_{new}$ consist of non-overlapping intervals, each containing at least one point. Therefore, given a batch size of $n$, we have $|\mathcal{I}|\leq n$ and $|\mathcal{G}_{new}|\leq n$. After receiving the gradients from clients, the server then computes $N$ observations $\mathbf{g}_{i}$ of size $d$, which requires $\mathcal{O}(Nd)$ operations, and adds them to the current set of strips $\mathcal{G}$ (lines 9 and 10). After this step, the size of $\mathcal{G}$ becomes at most $n+N$. To prepare the next-round attack (Alg. 2), the server updates the search space over the strips, which involves sorting the set $\mathcal{G}$ and iterating over it to compare vectors $\mathbf{g}_{i}$. This step incurs a computational cost of $\mathcal{O}(\max((N+n)\log(n+N),(n+N)d))$.
We then analyze the number of rounds $T$ required to isolate all inputs (i.e., achieve full batch recovery) with high probability. Assuming that the hyperplane $\mathbf{w}$ is drawn from a standard normal distribution and that $N\geq n$, we show that $T$ is $\mathcal{O}(\log\frac{n}{N})$ when the threshold $\epsilon$ is set to $\mathcal{O}(1/n^{2})$ (Appendix. D).
The final reconstruction step requires an additional $\mathcal{O}(nd)$ operations (line 12). With hardware optimized for parallel computation, the dimensional dependency on $d$ in the above steps can be effectively mitigated, resulting in a total complexity of $\mathcal{O}\left(\log\frac{n}{N}((N+n)\log(n+N))\right)$.
@@ 5 Experiments
We conduct experiments to demonstrate the efficacy of our proposed attack on both image and tabular data  [^3] Code: https://github.com/francescodiana99/cutting-through-privacy. Our evaluation encompasses two distinct datasets, each representing a different data modality. For image classification tasks, we use the ImageNet ILSVRC 2012 dataset [Russakovsky et al., 2015], that comprises a total of 1000 classes. As preprocessing steps, all images are rescaled to the $[0,1]$ range and resized to $224\times 224$ pixels. To evaluate performance on tabular data, we test the Human Activity Recognition Using Smartphones dataset [Anguita et al., 2013] (HARUS), which includes recordings of 30 subjects performing various activities while carrying a waist-mounted smartphone equipped with embedded sensors. The dataset comprises 561 features and 6 classification labels, representing the activity performed by each user. All features are scaled between 1-1 and $1$. Each client trains an FC-NN with two layers. Unless stated otherwise, the first layer consists of $N=1000$ neurons. For all the experiments, we assume that training is performed using FedSGD with full-batch updates. An image is considered reconstructed if it has SSIM $\geq 0.99$ relative to the true image (Figures 9 and 10 provide examples of reconstructed images). For tabular data, we consider an input to be perfectly reconstructed if the L2-norm difference between the true and recovered data point is <0.1<0.1.
We compare our method to the Curious Abandon Honesty (CAH) attack [Boenisch et al., 2023], allowing, for a fair evaluation, their attack to be performed across multiple communication rounds. Each experiment is repeated three times with different random seeds to ensure consistency and reliability of the results. Details on the configuration of both attacks can be found in App. E.
@@ 5.1 Experimental Results
Figure 3 illustrates the performance difference between our attack and CAH on the HARUS dataset. Two key trends emerge from our results. First, the performance gap increases with both the number of communication rounds and the batch size. This is clearly visible in the figure, with the largest performance difference ($97.75$ percentage points) observed at 50 rounds and a batch size of 4096. Second, different regions of the heatmap reveal the relative strengths of each approach. The top-left region (small batches, many rounds) suggests CAH’s effectiveness on smaller batches; the bottom row highlights its advantage in one-shot attacks. However, this advantage relies on their trap weights initialization, which requires the attacker to have access to some representative data to tune some hyperparameters (see App. E). This relative advantage disappears as the batch size increases. The top-right region (large batches, many rounds) highlights our attack’s ability to scale, demonstrating its potential for larger batches. This supports our claim that a sufficient number of rounds enables recovery for arbitrary large batches.
We also evaluate the effect of the number of neurons $N$ in the fully connected attack layer and its relationship with attack accuracy, see Table 1. While we observe that the percentage of reconstructed data increases with a larger number of neurons for both attacks, our attack outperforms the baseline in all but one combination of rounds and neuron counts. In many cases, our attack reconstructs nearly 100 times more inputs than CAH. We highlight a key distinction: the baseline reconstructs around 3.6% of the input after 50 rounds with 2000 neurons, i.e., using 100 000 different random hyperplanes. On the other hand, our attack manages to reconstruct 38% of the batch after only 10 rounds with only 500 neurons, i.e., using only 5000 different hyperplanes. In comparison, our attack reconstructs more than 10 times as many input points using 20 times fewer hyperplanes than the baseline, thanks to its efficient search procedure for correctly positioning the hyperplanes that isolate the inputs.
In terms of computational overhead, we evaluate our attack on the most computationally demanding setting—ImageNet with $n=1024$. In this case, the attack requires less than two seconds per round (primarily for lines 9 and 11 in Alg. 1) and under four seconds for the final image reconstruction phase (line 12), using an NVIDIA GeForce RTX 2080 Ti GPU. Since attacks are carried out across consecutive rounds, this additional overhead of approximately two seconds per round is negligible compared to the tens of seconds typically required to transmit the model in a single communication round.
We provide additional experimental results on HARUS, ImageNet, and CIFAR-10 [Krizhevsky, 2009] in App. F, including tests on a CNN architecture and extension to the multiple local steps setting, relaxing the assumption of full-batch updates. We also evaluate the effect of data heterogeneity and the effect of noise. These results confirm the same qualitative trends observed in our main results, further validating our conclusions.
@@ 6 Discussion
Secure aggregation [Bonawitz et al., 2017] is often proposed as a defense mechanism against privacy attacks, as it aggregates updates from multiple clients without revealing individual client updates to the server. However, the server still obtains the average of the gradients computed over the entire union of client datasets. With a sufficient number of communication rounds, our attack could still recover all underlying data, though not necessarily associate it with a specific client.
Therefore, the only robust defense against our attack that we are aware of would involve implementing local differential privacy (LDP) [Dwork, 2006, Bhowmick et al., 2019] on the client side. By adding noise to updates, LDP alters the coefficients associated to each input in (3), thus hindering precise localization and subsequent reconstruction of the data points, at the cost of decreased model utility. We conducted a preliminary evaluation of our attack’s performance when clients add Gaussian noise to their updates. Results in App. F show that noise perturbation decreases the proportion of inputs that our attack can perfectly recover. Nevertheless, our method remains effective: even when clients add noise, data reconstruction is still feasible and we consistently outperform the baseline in terms of reconstruction accuracy.
@@ 7 Future Works and Conclusion
Exploring the applicability of our attack to other machine learning tasks and architectures remains a promising direction for future work. Although our current approach is broadly applicable—by first recovering embeddings from the first fully connected layer and then applying a model inversion attack, following the methodology of Kariyappa et al. [2023] and Zhang et al. [2023]—the resulting reconstructions would still not be perfect. Therefore, developing analytical attacks that enable accurate, end-to-end reconstruction from more complex architectures and from different tasks remains an important and promising topic for future research.
Extending this attack to settings where clients perform multiple local epochs presents also an interesting challenge. In App. F, we carried out a preliminary evaluation by considering a scenario with one local epoch using minibatch stochastic gradient descent. However, effectively handling cumulative minibatch updates over multiple local epochs remains an open question.
Acknowledgements.
This research was supported in part by the French government, through the 3IA Côte d’Azur Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-19-P3IA-0002 and in part by the European Network of Excellence dAIEDGE under Grant Agreement Nr. 101120726. It was also funded in part by the Groupe La Poste, sponsor of the Inria Foundation, in the framework of the FedMalin Inria Challenge and by the EU HORIZON MSCA 2023 DN project FINALITY (G.A. 101168816). André Nusser was supported by the French government through the France 2030 investment plan managed by the National Research Agency (ANR), as part of the Initiative of Excellence of Université Côte d’Azur under reference number ANR-15-IDEX-01. The authors are grateful to the OPAL infrastructure from Université Côte d’Azur for providing resources and support.
@@ A Hyperplane-Based Data Reconstruction Attack in Federated Learning
By definition of the attack method, a point can only be reconstructed using (2) if it is linearly separated by injecting malicious weights. Thus, the number of recovered points is upper-bounded by the number of linearly separable points in the batch $\mathcal{B}_{{\mathbf{x}}}$, which corresponds to the number of vertices on its convex hull. This value is bounded by:
@@
    $O(n^{(d-1)/(d+1)})$ if $\mathcal{B}_{{\mathbf{x}}}$ is sampled uniformly at random from a unit ball [Raynaud, 1970], or
@@
    $O(\log^{d-1}n)$ if $\mathcal{B}_{{\mathbf{x}}}$ is sampled uniformly at random from the unit hypercube [Bentley et al., 1978], or
@@
    $O(\log^{(d-1)/2}n)$ if $\mathcal{B}_{{\mathbf{x}}}$ is sampled from a normal distribution with covariance matrix $\mathbf{I}_{d}$ [Raynaud, 1970].∎
@@ Appendix B Partial derivative with respect to the bias
$\frac{\partial{\mathcal{L}_{j}}}{\partial{b^{(1)}_{i}}}=\frac{\partial{\mathcal{L}_{j}}}{\partial{z^{(1)}_{i}}}=\frac{\partial{\mathcal{L}_{j}}}{\partial{{\mathbf{z}}^{(2)}}}\frac{\partial{{\mathbf{z}}^{(2)}}}{\partial{z_{i}^{(1)}}}.$ (11)